# Mask Image Watermarking

### v2

@@ Abstract
We present MaskMark, a simple, efficient, and flexible framework for image watermarking. MaskMark has two variants: (1) MaskMark-D, which supports global watermark embedding, watermark localization, and local watermark extraction for applications such as tamper detection; (2) MaskMark-ED, which focuses on local watermark embedding and extraction, offering enhanced robustness in small regions to support fine-grined image protection. MaskMark-D builds on the classical encoder-distortion layer-decoder training paradigm. In MaskMark-D, we introduce a simple masking mechanism during the decoding stage that enables both global and local watermark extraction. During training, the decoder is guided by various types of masks applied to watermarked images before extraction, helping it learn to localize watermarks and extract them from the corresponding local areas. MaskMark-ED extends this design by incorporating the mask into the encoding stage as well, guiding the encoder to embed the watermark in designated local regions, which improves robustness under regional attacks. Extensive experiments show that MaskMark achieves state-of-the-art performance in global and local watermark extraction, watermark localization, and multi-watermark embedding. It outperforms all existing baselines, including the recent leading model WAM for local watermarking, while preserving high visual quality of the watermarked images. In addition, MaskMark is highly efficient and adaptable. It requires only 20 hours of training on a single A6000 GPU, achieving 15× computational efficiency compared to WAM. By simply adjusting the distortion layer, MaskMark can be quickly fine-tuned to meet varying robustness requirements.
@@ 1 Introduction
Image watermarking [14] is a crucial technique for embedding imperceptible information into images, serving purposes such as copyright protection, content authentication, and provenance tracking. With the proliferation of AI-generated content (AIGC) [16; 18], the boundary between real and synthetic images has become increasingly blurred, making it especially important to develop robust watermarking schemes for content verification and traceability.
Traditional deep image watermarking methods [30; 21; 9] typically perform global watermark embedding and extraction, treating the entire image as a uniform entity. However, these global approaches suffer from several critical limitations. 1. Lack of local watermark extraction: When an image undergoes heavy tampering, such as inpainting [24; 26], the watermark may survive only in a small, local region that remains untouched by manipulation. In such cases, global methods often fail to extract the watermark effectively. 2. Inability to localize the watermark: Even if a watermark is successfully extracted from the image, global methods cannot localize which region actually contains the watermark, making it difficult for fine-grained forensic analysis and fair judgment. 3. Lack of local watermark embedding: In scenarios where only specific regions of an image are valuable and need protection, or when different parts of the image originate from different sources and require distinct watermarking, global embedding is inherently incapable of providing the flexibility and granularity.
@@
Based on the above analysis, we propose MaskMark, a simple, efficient, and flexible image watermarking framework. MaskMark introduces a masking mechanism during training to guide the model learn spatially aware embedding and extraction of local watermark signals. Depending on the stage which the mask is introduced, MaskMark offers two variants: MaskMark-D introduces the mask only during the decoding phase, enabling global watermark embedding while supporting local extraction. Specifically, by applying masks to retain only the selected regions of the watermarked images during extraction, the decoder is guided to identify which regions contain watermark signals and to focus on them for effective local extraction. MaskMark-ED introduces the mask during the encoding and decoding phases, allowing the embedding and extraction of the local watermark. In this setting, the encoder is trained to embed not only the watermark bits but also the spatial mask into the image. This allows the encoder to leverage the mask to adaptively allocate watermark strength to designated regions, while keeping the rest of the image nearly unaffected.
Extensive experiments demonstrate that MaskMark significantly outperforms existing baselines in both global and local watermark extraction, watermark localization, and local watermark embedding, while preserving image quality. Specifically, for local watermark extraction, MaskMark achieves a nearly 100% extraction accuracy even when only 5% of the image carries watermark signals. In terms of watermark localization, MaskMark demonstrates high precision in identifying watermark regions. Furthermore, although not explicitly trained for multi-watermark embedding, MaskMark maintains strong extraction and localization performance even when embedding up to 5 distinct watermarks in a single image. More importantly, MaskMark exhibits strong robustness across a wide range of distortions, including geometric distortions that typically break many existing watermarking methods. In addition to its effectiveness, MaskMark is highly efficient. Training the encoder-decoder model requires only approximately 20 hours on a single A6000 GPU, which is 15× less compute than the recent state-of-the-art local watermarking model WAM [19]. MaskMark also scales effortlessly to different bit lengths (e.g., 32, 64, and 128), whereas WAM is inherently limited to 32-bit embedding and does not scale beyond that. Furthermore, MaskMark supports fast fine-tuning for different use cases. For example, MaskMark can reach a near 100% extraction accuracy against VAE-based adaptive attacks after just 20k training steps. These advantages make MaskMark a practical, efficient, and scalable solution for real-world applications.
@@ 2.1 Image Watermarking
Image watermarking techniques can generally be categorized into two types: global watermarking and local watermarking methods. Global watermarking methods aim to extract watermark information from the entire image. Most traditional deep learning-based approaches fall into this category. These methods focus on achieving robustness against various types of distortions, ensuring that the embedded watermark can still be reliably recovered even when the image undergoes degradation. For example, MBRS [9] specifically targets robustness against JPEG compression. Methods like StegaStamp [21] and PIMoG [4] are designed to handle real-world physical distortions such as screen-shooting and print-shooting. More recent approaches like ZoDiac [25] and SuperMark [5] tackle adaptive attacks, while Robust-Wide [7] and VINE [13] focus on robustness against instruction-driven image editing.
In contrast, local watermarking methods focus on extracting watermark information from a specific region of the image. Recent methods, such as WAM [19] and our proposed MaskMark, belong to this category. WAM treats watermark extraction as a segmentation task [11], predicting watermark bits at the pixel level and then averaging these per-pixel predictions to obtain the final result. While this fine-grained approach enables local watermark extraction, it also presents several challenges. 1. Limited extraction from small regions: When the watermarked area is very small, only a few pixels contribute to extraction, making naive averaging unreliable, especially under distortions. 2. Lack of scalability with longer messages: WAM struggles to scale beyond 32-bit messages, as training becomes increasingly difficult with longer bit lengths. 3. High computational cost: Training WAM is resource-intensive, requiring eight V100 GPUs for nearly a week, which limits its practicality. 4. Lack of native local embedding: WAM embeds watermarks globally and then crops for local focus, introducing embedding losses that reduce extraction robustness.
@@ 2.2 Watermark Localization
Watermark localization [28; 6; 19] refers to the ability to determine which regions of a watermarked image still contain watermark information after modifications. This capability enables the identification of unaltered content, serving as an active detection mechanism for tamper localization. Currently, image watermarking techniques primarily adopt two paradigms for watermark localization. The first paradigm embeds a one-dimensional copyright watermark alongside a two-dimensional localization watermark in the original image. During extraction, localization is based on the fragility of the localization watermark, which cannot be fully recovered from a modified image. Key methods in this category include EditGuard [28] and OmniGuard [29]. EditGuard embeds a solid-color template within the host image and attempts to recover it from a modified version. The difference between the recovered and original templates is calculated at each pixel, and a threshold-based decision identifies watermark-preserved regions. OmniGuard improves upon EditGuard by embedding a natural image as the template, enhancing fidelity. It also introduces a Degradation-aware Tamper Extractor, improving robustness in detecting tampered regions under distortion. This paradigm requires parallel extraction of both copyright and localization watermarks, which may affect image quality. Moreover, both watermarks need independent robustness, and the presence of the template watermark does not guarantee the presence of the copyright watermark.
The second paradigm, in contrast, embeds only a one-dimensional copyright bitstream and directly determines the presence or absence of watermark information at each pixel to achieve localization. Methods such as WAM [19] and our proposed MaskMark fall under this category. WAM employs a decoder that simultaneously performs pixel-wise watermark presence detection and copyright bit extraction. In contrast, our MaskMark incorporates a dedicated localization module within the decoder, focusing solely on watermark presence detection at each pixel. This approach is more lightweight and easier to optimize. Compared to the first watermark localization paradigm, this method ensures that the local watermarked regions strictly correspond to the areas containing copyright watermark information, enhancing interpretability. Additionally, it guarantees both the robustness of copyright watermark extraction and the robustness of localization.
@@ 3.1 Design Principles
To address these challenges, we propose MaskMark-D, which introduces a basic mask mechanism during the decoding stage to guide the decoder in identifying and focusing on watermark-containing regions. To solve the first issue, we retain the watermark only in the masked regions and set other regions’ pixels to zero, training the decoder to extract watermarks from partially watermarked images. For the second issue, we replace the non-masked regions with the original clean image and add a watermark localization module in the decoder to differentiate between watermarked and non-watermarked areas, reducing interference from irrelevant content. These two strategies directly address the first two issues by enabling the decoder to extract watermarks from local regions, while also indirectly mitigating the third issue by guiding the encoder to evenly distribute the watermark under the end-to-end training, thus facilitating the decoder’s extraction process.
To further enhance the encoder’s ability to address the third challenge, we propose MaskMark-ED, which incorporates the mask during the encoding stage to explicitly guide watermark placement. In MaskMark-ED, the mask is embedded into the image along with the watermark bits during training. This enables the encoder to learn to actively concentrate the watermark within the selected regions based on the embedded mask, thereby further improving the robustness of local watermarking.
@@ 3.2 Training
Figure 1: The overall end-to-end training pipeline of MaskMark. (1) In the Mask Generation stage, we generate candidate masks from four predefined types and randomly select one mask $M$ for the subsequent stages. (2) In the Watermark Embedding stage, the encoder $\mathcal{E}$ embeds the watermark bits $W_{gt}$ into the original image $I_{orig}$, optionally using the mask $M$ (for MaskMark-ED), to produce the watermarked image $I_{wm}$. (3) In the Watermark Masking stage, the mask $M$ is used to fuse $I_{orig}$ and $I_{wm}$ (see Eq. 2), resulting in the fused image $I_{fuse}$, which is then subjected to a randomly selected distortion from a predefined distortion pool, yielding the distorted image $I^{{}^{\prime}}_{fuse}$. The masked region is then cropped using $M$ to obtain $I_{mask}$. (4) In the Watermark Extraction stage, the decoder $\mathcal{D}$ extracts the predicted mask $M_{pd}$ from $I^{{}^{\prime}}_{fuse}$ and the predicted watermark bits $W_{pd}$ from $I_{mask}$.
@@
The mask generation process constructs a pool of candidate masks with diverse types, from which one mask $M$ is randomly selected for each image during training. We follow a similar mask generation strategy as LaMa [20] for image inpainting, utilizing four distinct types of masks: Full Mask, Rectangle Mask, Irregular Mask, and Segment Mask. These masks can enhance the model’s ability to handle watermark embedding and extraction under diverse conditions from different perspectives: Full Mask enables the model to embed and extract watermarks across the entire image, serving as a fundamental capability. Rectangle Mask focuses on regularly shaped local regions, encouraging the model to operate within confined areas of varying sizes. Irregular Mask introduces complex, arbitrarily shaped regions to improve robustness in non-uniform contexts. Lastly, Segment Mask targets semantically meaningful areas by leveraging object masks from the MS-COCO dataset [12], helping the model generalize to real-world scenarios.
@@
To embed the watermark, we concatenate the original image $I_{orig}$ with the watermark feature $f$ along the channel dimension. For the MaskMark-D, this results in a tensor of shape $(3+C_{f},H,W)$, promoting global watermark embedding. For the MaskMark-ED, the optional mask $M$ is further concatenated, yielding a tensor of shape $(3+C_{f}+1,H,W)$. The mask guides the model to focus on the selected regions during training, enabling local embedding within those specified areas.
The concatenated tensor is then passed through a U-Net [11] to generate an intermediate encoded image $I_{enc}$. To obtain the final watermarked image $I_{wm}$, we apply a Just-Noticeable-Difference (JND) module [22], which modulates the embedding signal based on human visual sensitivity to enhance the perceptual quality:
$I_{wm}=I_{orig}+\mu\times\mathrm{JND}(I_{enc}-I_{orig}),$ (1)
@@
We describe how the decoder $\mathcal{D}$ extracts the predicted mask from $I^{\prime}_{fuse}$ and recovers the watermark bits from $I_{mask}$. To achieve these two objectives, $\mathcal{D}$ consists of two dedicated modules: a U2-Net [15] for mask prediction, and a U-Net [17] followed by a CNN for watermark extraction. Specifically, the U2-Net takes $I^{\prime}_{fuse}$ as input and predicts a mask $M_{pd}$. Meanwhile, the U-Net processes $I_{mask}$ to produce an intermediate feature $f^{\prime}$ with shape $(C_{f},H,W)$, which is then passed through a CNN to obtain the predicted watermark bits $W_{pd}$. Unlike the CNN used in the encoder $\mathcal{E}$, the CNN in the decoder $\mathcal{D}$ first applies several Conv–Norm–ReLU layers to further extract features from $f^{\prime}$. The resulting features are then interpolated to a fixed shape of $(1,l,l)$ and subsequently transformed by linear layers into a bit sequence of length $l$, yielding the final watermark prediction $W_{pd}$.

### v3

@@ Abstract
We present MaskWM, a simple, efficient, and flexible framework for image watermarking. MaskWM has two variants: (1) MaskWM-D, which supports global watermark embedding, watermark localization, and local watermark extraction for applications such as tamper detection; (2) MaskWM-ED, which focuses on local watermark embedding and extraction, offering enhanced robustness in small regions to support fine-grined image protection. MaskWM-D builds on the classical encoder-distortion layer-decoder training paradigm. In MaskWM-D, we introduce a simple masking mechanism during the decoding stage that enables both global and local watermark extraction. During training, the decoder is guided by various types of masks applied to watermarked images before extraction, helping it learn to localize watermarks and extract them from the corresponding local areas. MaskWM-ED extends this design by incorporating the mask into the encoding stage as well, guiding the encoder to embed the watermark in designated local regions, which improves robustness under regional attacks. Extensive experiments show that MaskWM achieves state-of-the-art performance in global and local watermark extraction, watermark localization, and multi-watermark embedding. It outperforms all existing baselines, including the recent leading model WAM for local watermarking, while preserving high visual quality of the watermarked images. In addition, MaskWM is highly efficient and adaptable. It requires only 20 hours of training on a single A6000 GPU, achieving 15× computational efficiency compared to WAM. By simply adjusting the distortion layer, MaskWM can be quickly fine-tuned to meet varying robustness requirements.
@@ 1 Introduction
Image watermarking (potdar2005survey, ) is a crucial technique for embedding imperceptible information into images, serving purposes such as copyright protection, content authentication, and provenance tracking. With the proliferation of AI-generated content (AIGC) (rombach2022high, ; saharia2022photorealistic, ), the boundary between real and synthetic images has become increasingly blurred, making it especially important to develop robust watermarking schemes for content verification and traceability.
Traditional deep image watermarking methods (zhu2018hidden, ; tancik2020stegastamp, ; jia2021mbrs, ) typically perform global watermark embedding and extraction, treating the entire image as a uniform entity. However, these global approaches suffer from several critical limitations. 1. Lack of local watermark extraction: When an image undergoes heavy tampering, such as inpainting (yu2023inpaint, ; zhang2023adding, ), the watermark may survive only in a small, local region that remains untouched by manipulation. In such cases, global methods often fail to extract the watermark effectively. 2. Inability to localize the watermark: Even if a watermark is successfully extracted from the image, global methods cannot localize which region actually contains the watermark, making it difficult for fine-grained forensic analysis and fair judgment. 3. Lack of local watermark embedding: In scenarios where only specific regions of an image are valuable and need protection, or when different parts of the image originate from different sources and require distinct watermarking, global embedding is inherently incapable of providing the flexibility and granularity.
@@
Based on the above analysis, we propose MaskWM, a simple, efficient, and flexible image watermarking framework. MaskWM introduces a masking mechanism during training to guide the model learn spatially aware embedding and extraction of local watermark signals. Depending on the stage which the mask is introduced, MaskWM offers two variants: MaskWM-D introduces the mask only during the decoding phase, enabling global watermark embedding while supporting local extraction. Specifically, by applying masks to retain only the selected regions of the watermarked images during extraction, the decoder is guided to identify which regions contain watermark signals and to focus on them for effective local extraction. MaskWM-ED introduces the mask during the encoding and decoding phases, allowing the embedding and extraction of the local watermark. In this setting, the encoder is trained to embed not only the watermark bits but also the spatial mask into the image. This allows the encoder to leverage the mask to adaptively allocate watermark strength to designated regions, while keeping the rest of the image nearly unaffected.
Extensive experiments demonstrate that MaskWM significantly outperforms existing baselines in both global and local watermark extraction, watermark localization, and local watermark embedding, while preserving image quality. Specifically, for local watermark extraction, MaskWM achieves a nearly 100% extraction accuracy even when only 5% of the image carries watermark signals. In terms of watermark localization, MaskWM demonstrates high precision in identifying watermark regions. Furthermore, although not explicitly trained for multi-watermark embedding, MaskWM maintains strong extraction and localization performance even when embedding up to 5 distinct watermarks in a single image. More importantly, MaskWM exhibits strong robustness across a wide range of distortions, including geometric distortions that typically break many existing watermarking methods. In addition to its effectiveness, MaskWM is highly efficient. Training the encoder-decoder model requires only approximately 20 hours on a single A6000 GPU, which is 15× less compute than the recent state-of-the-art local watermarking model WAM (sander2025watermark, ). MaskWM also scales effortlessly to different bit lengths (e.g., 32, 64, and 128), whereas WAM is inherently limited to 32-bit embedding and does not scale beyond that. Furthermore, MaskWM supports fast fine-tuning for different use cases. For example, MaskWM can reach a near 100% extraction accuracy against VAE-based adaptive attacks after just 20k training steps. These advantages make MaskWM a practical, efficient, and scalable solution for real-world applications.
@@ 2.1 Image Watermarking
Image watermarking techniques can generally be categorized into two types: global watermarking and local watermarking methods. Global watermarking methods aim to extract watermark information from the entire image. Most traditional deep learning-based approaches fall into this category. These methods focus on achieving robustness against various types of distortions, ensuring that the embedded watermark can still be reliably recovered even when the image undergoes degradation. For example, MBRS (jia2021mbrs, ) specifically targets robustness against JPEG compression. Methods like StegaStamp (tancik2020stegastamp, ) and PIMoG (fang2022pimog, ) are designed to handle real-world physical distortions such as screen-shooting and print-shooting. More recent approaches like ZoDiac (zhang2024ZoDiac, ) and SuperMark (hu2024supermark, ) tackle adaptive attacks, while Robust-Wide (hu2024robust, ) and VINE (lu2024robust, ) focus on robustness against instruction-driven image editing.
In contrast, local watermarking methods focus on extracting watermark information from a specific region of the image. Recent methods, such as WAM (sander2025watermark, ) and our proposed MaskWM, belong to this category. WAM treats watermark extraction as a segmentation task (kirillov2023segment, ), predicting watermark bits at the pixel level and then averaging these per-pixel predictions to obtain the final result. While this fine-grained approach enables local watermark extraction, it also presents several challenges. 1. Limited extraction from small regions: When the watermarked area is very small, only a few pixels contribute to extraction, making naive averaging unreliable, especially under distortions. 2. Lack of scalability with longer messages: WAM struggles to scale beyond 32-bit messages, as training becomes increasingly difficult with longer bit lengths. 3. High computational cost: Training WAM is resource-intensive, requiring eight V100 GPUs for nearly a week, which limits its practicality. 4. Lack of native local embedding: WAM embeds watermarks globally and then crops for local focus, introducing embedding losses that reduce extraction robustness.
@@ 2.2 Watermark Localization
Watermark localization (zhang2024editguard, ; hu2025videoshield, ; sander2025watermark, ) refers to the ability to determine which regions of a watermarked image still contain watermark information after modifications. This capability enables the identification of unaltered content, serving as an active detection mechanism for tamper localization. Currently, image watermarking techniques primarily adopt two paradigms for watermark localization. The first paradigm embeds a one-dimensional copyright watermark alongside a two-dimensional localization watermark in the original image. During extraction, localization is based on the fragility of the localization watermark, which cannot be fully recovered from a modified image. Key methods in this category include EditGuard (zhang2024editguard, ) and OmniGuard (zhang2024omniguard, ). EditGuard embeds a solid-color template within the host image and attempts to recover it from a modified version. The difference between the recovered and original templates is calculated at each pixel, and a threshold-based decision identifies watermark-preserved regions. OmniGuard improves upon EditGuard by embedding a natural image as the template, enhancing fidelity. It also introduces a Degradation-aware Tamper Extractor, improving robustness in detecting tampered regions under distortion. This paradigm requires parallel extraction of both copyright and localization watermarks, which may affect image quality. Moreover, both watermarks need independent robustness, and the presence of the template watermark does not guarantee the presence of the copyright watermark.
The second paradigm, in contrast, embeds only a one-dimensional copyright bitstream and directly determines the presence or absence of watermark information at each pixel to achieve localization. Methods such as WAM (sander2025watermark, ) and our proposed MaskWM fall under this category. WAM employs a decoder that simultaneously performs pixel-wise watermark presence detection and copyright bit extraction. In contrast, our MaskWM incorporates a dedicated localization module within the decoder, focusing solely on watermark presence detection at each pixel. This approach is more lightweight and easier to optimize. Compared to the first watermark localization paradigm, this method ensures that the local watermarked regions strictly correspond to the areas containing copyright watermark information, enhancing interpretability. Additionally, it guarantees both the robustness of copyright watermark extraction and the robustness of localization.
@@ 3.1 Design Principles
To address these challenges, we propose MaskWM-D, which introduces a basic mask mechanism during the decoding stage to guide the decoder in identifying and focusing on watermark-containing regions. To solve the first issue, we retain the watermark only in the masked regions and set other regions’ pixels to zero, training the decoder to extract watermarks from partially watermarked images. For the second issue, we replace the non-masked regions with the original clean image and add a watermark localization module in the decoder to differentiate between watermarked and non-watermarked areas, reducing interference from irrelevant content. These two strategies directly address the first two issues by enabling the decoder to extract watermarks from local regions, while also indirectly mitigating the third issue by guiding the encoder to evenly distribute the watermark under the end-to-end training, thus facilitating the decoder’s extraction process.
To further enhance the encoder’s ability to address the third challenge, we propose MaskWM-ED, which incorporates the mask during the encoding stage to explicitly guide watermark placement. In MaskWM-ED, the mask is embedded into the image along with the watermark bits during training. This enables the encoder to learn to actively concentrate the watermark within the selected regions based on the embedded mask, thereby further improving the robustness of local watermarking.
@@ 3.2 Training
Figure 1: The overall end-to-end training pipeline of MaskWM. (1) In the Mask Generation stage, we generate candidate masks from four predefined types and randomly select one mask $M$ for the subsequent stages. (2) In the Watermark Embedding stage, the encoder $\mathcal{E}$ embeds the watermark bits $W_{gt}$ into the original image $I_{orig}$, optionally using the mask $M$ (for MaskWM-ED), to produce the watermarked image $I_{wm}$. (3) In the Watermark Masking stage, the mask $M$ is used to fuse $I_{orig}$ and $I_{wm}$ (see Eq. 2), resulting in the fused image $I_{fuse}$, which is then subjected to a randomly selected distortion from a predefined distortion pool, yielding the distorted image $I^{{}^{\prime}}_{fuse}$. The masked region is then cropped using $M$ to obtain $I_{mask}$. (4) In the Watermark Extraction stage, the decoder $\mathcal{D}$ extracts the predicted mask $M_{pd}$ from $I^{{}^{\prime}}_{fuse}$ and the predicted watermark bits $W_{pd}$ from $I_{mask}$.
@@
The mask generation process constructs a pool of candidate masks with diverse types, from which one mask $M$ is randomly selected for each image during training. We follow a similar mask generation strategy as LaMa (suvorov2022resolution, ) for image inpainting, utilizing four distinct types of masks: Full Mask, Rectangle Mask, Irregular Mask, and Segment Mask. These masks can enhance the model’s ability to handle watermark embedding and extraction under diverse conditions from different perspectives: Full Mask enables the model to embed and extract watermarks across the entire image, serving as a fundamental capability. Rectangle Mask focuses on regularly shaped local regions, encouraging the model to operate within confined areas of varying sizes. Irregular Mask introduces complex, arbitrarily shaped regions to improve robustness in non-uniform contexts. Lastly, Segment Mask targets semantically meaningful areas by leveraging object masks from the MS-COCO dataset (ms-coco, ), helping the model generalize to real-world scenarios.
@@
To embed the watermark, we concatenate the original image $I_{orig}$ with the watermark feature $f$ along the channel dimension. For the MaskWM-D, this results in a tensor of shape $(3+C_{f},H,W)$, promoting global watermark embedding. For the MaskWM-ED, the optional mask $M$ is further concatenated, yielding a tensor of shape $(3+C_{f}+1,H,W)$. The mask guides the model to focus on the selected regions during training, enabling local embedding within those specified areas.
The concatenated tensor is then passed through a U-Net (kirillov2023segment, ) to generate an intermediate encoded image $I_{enc}$. To obtain the final watermarked image $I_{wm}$, we apply a Just-Noticeable-Difference (JND) module (wu2017enhanced, ), which modulates the embedding signal based on human visual sensitivity to enhance the perceptual quality:
$I_{wm}=I_{orig}+\mu\times\mathrm{JND}(I_{orig})\times(I_{enc}-I_{orig}),$ (1)
@@
We describe how the decoder $\mathcal{D}$ extracts the predicted mask from $I^{\prime}_{fuse}$ and recovers the watermark bits from $I_{mask}$. To achieve these two objectives, $\mathcal{D}$ consists of two dedicated modules: a U2-Net (qin2020u2, ) for mask prediction, and a U-Net (ronneberger2015u, ) followed by a CNN for watermark extraction. Specifically, the U2-Net takes $I^{\prime}_{fuse}$ as input and predicts a mask $M_{pd}$. Meanwhile, the U-Net processes $I_{mask}$ to produce an intermediate feature $f^{\prime}$ with shape $(C_{f},H,W)$, which is then passed through a CNN to obtain the predicted watermark bits $W_{pd}$. Unlike the CNN used in the encoder $\mathcal{E}$, the CNN in the decoder $\mathcal{D}$ first applies several Conv–Norm–ReLU layers to further extract features from $f^{\prime}$. The resulting features are then interpolated to a fixed shape of $(1,l,l)$ and subsequently transformed by linear layers into a bit sequence of length $l$, yielding the final watermark prediction $W_{pd}$.